Do not just fill the blanks. Please explain how you get each blank. show your work. Thanks

1. Complete the One-Way ANOVA summary table (you do NOT need raw data to do this) df MS Fobs CV Source 40 Between (group)80 Within/Error Total 100 14
Answer #1

Sum of squares (SS):  

SSwithin = SSTotal - SSBetween = 100 - 80 = 20

SSwithin = 20

Dgrees of Freedom (DF) :

DF_{Between} = \frac {SS_{Between}} { MS_{Between}} = \frac {80}{40} = 2

DFBetween = 2

DFWithin = DFTotal - DFBetween = 14 - 2 = 12

Mean sum of squares:

MS_{within} = \frac {SS_{within}} { DF_{within}} = \frac {20}{12} = 1.6667

MSwithin = 1.67

F_{obs} = \frac {MS_{Betwewn}} { MS_{within}} = \frac {40}{1.67} = 23.9521

Fobs = 23.95

ANOVA summary table:

Source SS df MS Fobs
Between(group) 80 2 40 23.95
Within/Error 20 12 1.67
Total 100 14